OpenStudy (anonymous):

what is the gradient of this line y=2x+4

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

y = gradient(x) + intercept

OpenStudy (amistre64):

what is the gradient match up to?

OpenStudy (anonymous):

2???

OpenStudy (amistre64):

exactly :)

OpenStudy (anonymous):

okay i have another for u

OpenStudy (amistre64):

and it aint even my birtday; you shouldnt have

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

2y-2x=4

OpenStudy (amistre64):

lol... tricky

OpenStudy (amistre64):

its gonna be 1

OpenStudy (anonymous):

can u please explain how u got 1

OpenStudy (amistre64):

we can 'solve' for y: 2y -2x = 4 ; +2x 2y = 2x +4 ; /2 y = (2/2)x +(4/2) the gradient is 2/2 = 1